Circular No. 6076

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 SC, 1993 RO These two transneptunian objects were recovered by M. Kidger (on

Aug. 15 and 16 with the 2.5-m Isaac Newton reflector) and by D. Jewitt and J. Luu (on Sept. 5 and 7 with the 10-m Keck reflector and University of Hawaii 2.2-m reflector), respectively. General orbit solutions indicate a = 41.4 AU, e = 0.27 and a = 40.5 AU, e = 0.25, figures that tend to support the undersigned’s hypothesis (IAUC 5983, 5985) that orbital stability is maintained by the 2:3 mean-motion resonance with Neptune. The new observations, together with orbital elements that assume this hypothesis, are given on MPEC 1994-Q04 and 1994-R06. It is important to note that the validity of the hypothesis has not been proven, and further observations are required in order to demonstrate whether the orbits are in fact stable.

NOTICE REGARDING DESIGNATIONS AND NAMES OF COMETS

At a meeting on Aug. 24 the IAU General Assembly adopted a resolution that will result in a change in the cometary designation system effective 1995 Jan. 1. The present year/letter and year/Roman numeral systems will be replaced by a single system resembling that for minor planets, with objects recorded by the halfmonth. Thus the third comet reported as discovered during the second half of Feb. 1995 would be designated 1995 D3. Much of the rationale for the change stems from the difficulty that is sometimes apparent in deciding whether a particular object is in fact a comet or a minor planet (e.g., 2060 Chiron, 1990 UL_{3} = 1990p = 1990 XVI P/Shoemaker-Levy 2, 1977t = 1977 YA, etc.). When appropriate, the nature—or suggested nature—of an object can be indictated by preceding the designation with A/ (for minor planet), C/ (for comet), P/ (as now, for periodic comet), etc. In a process analogous to the permanent numbering of minor planets, sequential numbers will be defined for comets whose periodicity has been well established, and in such cases these numbers should immediately precede the P/ notation; “routine” recoveries of these periodic comets will not in the future receive additional designations. The system also allows the recognition of components of comets that have split, of cometary images that have been noted on the Palomar Sky Survey (for example), and the indication that some periodic comets have been lost (or, in one recent famous case, destroyed). The IAU resolution, which was published in its entirety on MPC 23803–23804 on Aug. 28, also touched briefly upon the matter of comet names. It affirmed the intent “to retain in general terms the tradition of naming comets for their discoverers” while at the same time noting the need “to ensure fairness and simplicity”.

1994 September 10 Brian G. Marsden